Current Status Summary

    Station Monitoring

    Monitoring of the station environment and the status of the LCU can be found at: https://lofarx.mpifr-bonn.mpg.de/cacti/graph_view.php?action=tree&tree_id=1

     

     

    Hardware Status as of 4. June 2013

    Electronics:

    All RCU modules plugged in.

    High Band Antennas:

    All tiles are working.

    In tile 54 (RCUs 108/109) the element 9 is not communicating.

    In tile 9 (RCUs 18,19) the element 16 shows strange features in the spectrum even though it has been recently replaced.

    Low Band Antennas:

    A number of LBAs show too little power, two show too high power (especially at low frequencies). This needs to be addressed at the next maintenance session.

    Some (even-numbered) dipoles show little to no power in the spectrum (like as if the LNA is not working). The problem is not always consistent: some seem to recover without active intervention, some seem to recover after pulling and re-connecting the RCU.

    Date RCUs with low power in rcumode 3
    Maintenance 16.10.2012 * * * * * * * * * * * * * * * * * * *
    Wed 1.02.2012         684 724   78       1044 1304 138 1482 150 1524   1784
    Wed 18.01.2012 * *   *   72   78         130 138 148 150 152   178
    Wed 21.9.2011 8 12 * 542 68 72   78 82 *     130 1381 148 150 1521 164 178
    Thu 11.8.2011 8 12 44 54   72   78 82 86     130 1381 148 150 1521 164 178
    Fri. 8.7.2011A 8 12 44   68 72 743 78 82 86     130   148 150 1522 164 178
    Mon 4.7.2011 8 12 44     72 741 78 822 86     130   148 150   164 178
    Fri 1.7.2011A 8 12 44   68 72   78   86     130   148 150   164 178
    Wed 29.6.2011A 8   44     72   78   86 92     138 148 150   164  

    Notes:
    A No attempt to recover signal by switching rcumodes done.
    1Recovered after switching to rcumode 1 and back (a number of times)
    2Showed changing power levels.
    3 At 50-50 MHz about 10-15 dB reduced at 20 MHz about 30 dB reduced.
    4 Normal power at some times, reduced power at others

    * Replaced during maintenance visit, or before

    Computer Status as of 4. June 2013:

    Data Taking PCs

    All lofarAN machines have new RAID disks configured ad RAID5 (12.7 TB storage space in total per machine).

    The lofar BN machines have their disks in a RAID0 configuration, with the system as part of the RAID0.

    Right now (4. June 2013) lofarB1 is offline due to a broken disk.

    Server

    lofarsrv is up

    List of spares as of 4. June 2013:

    Description Quantity (Note)
    spare disks (2TB) 7
    RCU - Modules

    8

    + 1 with broken LBH connector (used for demonstrations)

    HBA FE - Modules 0
    HBA Summators 1 (communication, without voltage regulator)
    "good, old style" LBA heads 8
    good, new-style LBA heads 0
    "LBA, HBA mechanics goodie bag" 1
    4 GByte TBB Memory modules 4 (spare, surpuls ???)
    2 GByte TBB Memory modules (old modules) 48 (DDR 2, 667, CL5, ECC, Reg)

     

    Important or Interesting Webpages

    Tag page (Edit tags)
    • No tags
    Page statistics
    14740 view(s), 63 edit(s) and 7862 character(s)

    Comments

    You must login to post a comment.

    Attach file

    Attachments